OpenStudy (rainbow_rocks03):

Find the probability of the event when a number cube with faces labeled 1– 6 is rolled once. a number greater than 4 A. 2/3 B. 1/3 C. 1/6 D. 1/2

OpenStudy (rainbow_rocks03):



OpenStudy (rainbow_rocks03):

@CallMeKiki @cutiemarklove @KendrickLamar2014 @LegendarySadist @Legends

OpenStudy (rainbow_rocks03):

Find the probability of the event when a number cube with faces labeled 1– 6 is rolled once. a number greater than 4

OpenStudy (anonymous):

i think is d

OpenStudy (rainbow_rocks03):

@LegendarySadist ?

OpenStudy (anonymous):

So there are 6 possibilities and you want the chance of getting 2 of them. So \[\frac{ 2 }{ 6 }\] is the chance. Just simplify for the answer.

OpenStudy (rainbow_rocks03):

How do I simplify 2/6 ?

OpenStudy (rainbow_rocks03):

never mind so the answer is A?

OpenStudy (anonymous):

Find the similar number, which is 2 in this case\[\frac{ 2 }{ 6 }\div \frac{ 2 }{ 2 }\to \frac{ 2/2 }{ 6/2 }=\frac{ 1 }{ 3 }\] So the answer is 1/3

OpenStudy (rainbow_rocks03):

oh..thx

OpenStudy (anonymous):

No problem :)